There are 3 blue balls, 5 red balls, and 4 white balls in a bag of balls. If a person selects two of the balls, what is the probability that the second one is blue given that the first one was white?
I have several problems that are similar to the above question. I have no idea where to start.
Thank you in advance.

Answer by rothauserc(4718)

There are 3 blue balls, 5 red balls, and 4 white balls in a bag of balls.
The probability(Pr) that the first ball drawn is white is 4 / 12 = 1/3,
now the Pr of drawing a blue ball is 3 / 11, therefore
The Pr of drawing a blue ball given the first is white = (1/3)*(3/11) = 3/33 = 1/11
Note that this answer assumes that the white ball is not put back in the bag.